OpenStudy (o.lee14):

Marcel converts the frequency table to a conditional relative frequency table by row. Which value should he use for Y? Round to the nearest hundredth. 0.19 0.45 0.82 0.90

OpenStudy (sterling234):

I would just plug each answer in for example y+z=1.0 right so one of these 0.19 0.45 0.82 0.90 answer choices are right lets go with the first one 0.19+z=1.0 subtract 0.19 from both sides= z= 0.81 this cant be true cause it also says x+z=0.57 and 0.81 is higher than that and x can't be a negative so choice a is out.

OpenStudy (sterling234):

I hope that makes sense lol